OpenStudy (anonymous):

Find the exact length of the curve. y = ln(sec x), 0 ≤ x ≤ π/4

OpenStudy (freckles):

\[\int\limits_{0}^{\frac{\pi}{4}}\sqrt{1+(y')^2} dx \]

OpenStudy (freckles):

do you recognize this formula?

OpenStudy (anonymous):

Yes! I'm just stuck trying to integrate \[\int\limits_{0}^{\pi/4} \sqrt{1+\tan^2} dx\]

OpenStudy (freckles):

what is 1+tan^2(x) think of an identity

OpenStudy (freckles):

Pythagorean identity

OpenStudy (anonymous):

these problems cooked to within an inch of their life think of all the effort it took to make a problem where you could take the derivative, square it, add 1, take the square root, and then come up with something integrable i tried it once, it took over half an hour maybe over an hour

OpenStudy (freckles):

lol That is why I give up and only gave problems from the book as test questions

OpenStudy (anonymous):

i know right? i was kind of proud of myself when i was done

OpenStudy (anonymous):

Oh sec^2(x)!

OpenStudy (freckles):

yeppie!

OpenStudy (freckles):

and sec is positive on the interval 0 to pi/4 so sqrt(sec^2(x))=sec(x)

OpenStudy (anonymous):

Thank you so much! :)

OpenStudy (freckles):

np